two cars leave town at the same time going in opposite directions. one car travels 44mi/h and the other travels 55mi/h. in how many hours will they be 297 miles apart?
Found 2 solutions by stanbon, rfer:
Answer by stanbon(57337)

two cars leave town at the same time going in opposite directions. one car travels 44mi/h and the other travels 55mi/h. in how many hours will they be 297 miles apart?
---
They are separating at the rate of 44+55 = 99 mph
---
time = distance/rate
tim = 297/99 = 3 hrs
================================
Cheers,
Stan H.

Answer by rfer(12657)

moving apaprt at 99 mph
297/99=3 hrs